[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Nov 23, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Wed Nov 23 16:00:45 UTC 2016

Here is today's summary of NHC computer operations:
--- Craig Mattocksi


1. The problem with the script that transfers early Wind Speed  
Probabilities from WCOSS to AWIPS cannot be addressed until TSB AWIPS  
SysAdmin Chris Mello returns after the Thanksgiving holiday (next Monday).

2. TSB Developer Matt Sardi added a menu option for "basin crossers" to the  
Danger Graphic application to correct the storm label problem and provided  
written instructions to TAFB forecasters through email.

3. When TAFB forecaster Patricia Wallace tried to reboot the cumulus  
workstation at the start of her shift yesterday, she inadvertantly hit the  
keyboard with the monitor. TSB Developer Matt Onderlinde had Patricia  
reboot the workstation. She was then able to employ the new Danger Graphic  
script for basin-crossing storms for TS Otto.

4. For the second day in a row, the NHC "product tracker"  
http://www2.nhc.noaa.gov/dataflow/frames.html showed the Offshore Waters  
Forecast for Southwest and Tropical Atlantic and Caribbean Sea  
(FZNT23-KNHC) as not being sent, even though it was resent and confirmed to  
be in AWIPS and on the NHC website. TSB will investigate.
1. The 12z GFS and EC model data arrived 30-45 minutes late yesterday. This  
has happened recently during upgrades of the NCEP Century Link network  
circuits and routers. TSB will monitor the network load using  
the "scrutinizer" tool.

2. When opening a terminal in the 'atlantic' account on ridges, the  
message "id: cannot find name for group ID 300" appeared at the top of each  
terminal window. NCO Linux SysAdmin Brian Maher fixed this problem.

3. The system clock on the Atlantic ATCF workstation was about 35 seconds  
ahead of the other system clocks on the ops floor. An NCO Help Desk ticket  
was submitted to make sure the workstation was synchronizing with an  
appropriate NTP time server. NCO Linux SysAdmin Brian Maher fixed this time  
drift problem.
Other items:
1. The NOAAPORT server (nport2.nhc.noaa.gov) stopped ingesting data last  
night, which meant that NHC was down to one functioning LDM feed (TOC Line  
162 socket feed server, logger.nhc.noaa.gov). TSB Developer Matt Sardi  
performed an LDM server stop/start and re-created the LDM queue, without  
success. An NCO Linux SysAdmin was needed to stop/start other services as  
root in order to get the data flowing again.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161123/bb9d0ef4/attachment.html 

More information about the Ncep.nhc.nco_contacts mailing list